You own a portfolio that has $2,450 invested in Stock A and $3,750 invested in Stock B. If the expected returns on these stocks are 10 percent and 15 percent, respectively, what is the expected return on the portfolio?(Do not round your intermediate calculations.)
Answer #1

Total investment=2,450+3,750=$6200

Portfolio return=Respective return*Respective weight

=(2,450/6200*10)+(3,750/6200*15)

=13.02%(Approx).